Deteremine the number of diagonals of a regular 79-gon.

Thank you

Answer by ewatrrr(24785)

 
Hi,
Deteremine the number of diagonals of a regular 79-gon.
Number of diagonals = n(n-3)2 = 79*76/2 = 3002